If a permutation of the word "white" is selected at random, find the probability that has the consonants and vowels alternating.
Answer by Edwin McCravy(20060)

They can only be in this form CVCVC, where C is a consonant 
and V is a vowel.

Choose a consonant to go first any of 3 ways, W,H, or T
That leaves 2 ways to choose a consonant to go third.
That leaves 1 way to choose a consonant to go fifth.
Choose a vowel to go second either of two ways, I or E.
That leaves 1 vowel to go fourth.

Answer 3*2*1*2*1 = 3!2! = 6*2 = 12 ways

 1. WIHET
 2. WEHIT
 3. WITEH
 4. WETIH
 5. HIWET
 6. HEWIT
 7. HITEW
 8. HETIW
 9. TIWEH
10. TEWIH
11. TIHEW
12. TEHIW

Edwin